0740: Day starts: Drank 1L of water to hydrate body and counteract dehydration caused by long term exposure to dry, cool air conditioned laboratory.
0745: System checks. Inspected the ion gauge pressures, the status of all pumps and LN2 level. Inserted MIG into beam path in preparation for BEP flux measurements.
0746: Outgas Ga. Executed a program in the Ga cell’s PID controller than ramps up to 975°C in 30 minutes, holds for 30 minutes and then ramps down to 950°C and holds. NB: The growth schedule for this week is more GaAs NW samples grown on Si(111) wafers. Only a single Ga cell is needed for this purpose.
0748: Preparation chamber LN2. Opened the preparation chamber’s LN2 solenoid and cooled the TSP pumps housing in preparation for the sample outgas. The manual fill was observed with a critical eye. The intention to implement “automatic LN2 fill of preparation chamber” was reaffirmed whilst shaking a fist at the heavens.
0750: Outgas sample. One of the two samples loaded last Friday evening was moved to the outgas stage in the preparation chamber. The outgas program was executed in PID controller: Ramp from 50°C to 400°C in 24 minutes, hold for 1 hour, ramp down to 100°C and hold.
0800: Update log. Sample locations were updated on the “locations” tab of the log spread sheet to reflect the new sample locations and status.
0810: As flux check. As shutter and valve were opened. Fluxes were taken at several valve positions consistent with previous days. Fluxes are recorded in the “flux check record” tab of the log spread sheet.
0850: As flux check completed. Gathered flux values were compared with legacy data from the log spread sheet for continuity checks. 40 minutes of the day were consumed simply checking the As flux, plus another few minutes spent daydreaming about an automatic flux checking software extension. The intention to implement “automatic flux checking” was reaffirmed whilst shaking fist at the heavens. As shutter is closed and background As4 is pumped from the system. Vacuum recovery progresses prior to Ga flux check.
0900: Ga flux check. Ga shutter was opened for 60 seconds and the flux was recorded, averaged over 10 seconds. The shutter was then closed, the background flux was sampled, subtracted and the “flux check record” tab of the log spread sheet was updated. Current flux value is compared to legacy data for continuity.
0910: RHEED. Retracted MIG. Turn RHEED gun on and set to standby value in preparation for growth. Inspected outgas program for sample 1: 45 minutes remain. Top up preparation chamber’s LN2.
0930: Confer with colleague. Discussed recent results with a colleague and outline strategy for samples to grow today.
1010: Sample 1: Start. Transferred sample from outgas stage to growth chamber and manually began recipe. Lament the current lack of recipe execution software. Attempt to transfer the second sample to the outgas stage failed due to misalignment of cassette inside FEL. Shake fist. Seal FEL, vent, dismount cassette. Used the opportunity to load a new sample and dismount the Ta outgas plate. Reload cassette and double check alignment. Pump down and degas FEL. Cost of misalignment: 4 hours. However loading the second sample means a total of 3 samples can be grown today if I work 12 hours…
1215: Sample 1: End. Sample 1 complete. Left sample to cool to 100°C in preparation for transfer.
1220: Lunch time… and consumption of another litre of water.
1315: Sample load/unload. Removed sample 1 from the growth chamber and place it in FEL. First sample from FEL was placed it in outgas stage. Second sample from FEL was placed in the storage cassette. Vented FEL, removed sample 1 and reloaded platen with new wafer. Pump down FEL and begin degas. Gave sample 1 to colleague for SEM analysis.
1500: Sample 2 start. With sample 2 (finally) outgassed, it was transferred to the growth chamber and the recipe started. Sample 3 was then transferred to the outgas stage and the standard outgas recipe was executed.
1515: Platen cleaning with wet and dry abrasive sheet. Abrasive sheet was too fine, retry tomorrow.
1700: Sample reshuffle. With the outgas of sample 3 complete, it was moved to the storage cassette. The newly degassed sample was moved from the FEL to the outgas stage in preparation for an outgas at the end of the day, and hence growth tomorrow.
1720: Sample 2 end. Sample 2 was left to cool for 30 minutes and was then transported into the FEL pending removal tomorrow morning. Sample 3 was transferred to the growth chamber. The outgas was performed on the new sample for tomorrow.
1750: Sample 3 start. The growth of sample 3 was started.
1800: Paperwork. Growth sheets were double checked. Run log was updated. All files were backed up to online server. Paper lab book was updated.
1930: Sample 3 end. After growth, sample 3 was transferred to the FEL pending removal (along with sample 2) tomorrow.
1940: Day ends. Drink another litre of water. Go home.
Total 12 hours. Growth: 3 samples. Samples/hour: 0.25.